OpenStudy (anonymous):

If f(x)=x4cosx, then f′(x)=

OpenStudy (campbell_st):

you will need the product rule if \[y = f(x) \times g(x)\] then \[\frac{dy}{dx} = f(x) \times g'(x) + g(x) \times f'(x)\] so let \[f(x) = x^4 ..... and ....g(x) = \cos(x)\] can you find the derivative of f(x) and g(x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

xcos(4) ?

OpenStudy (campbell_st):

umm do you know about the product rule. and is the inital function correct..?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

oops let \[g(x) = x^4\] and \[h(x) = \cos(x)\] what are the derivatives of g(x) and h(x)

OpenStudy (anonymous):

g'(x).h(x)+g(x).h'(x)

OpenStudy (campbell_st):

thats it... thats the product rule